$SPAD/src/input grdef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- test input for GraphicsDefaults package
--Operations to set graphic defaults:
adaptive == false
adaptive == true
clipPointsDefault(false)
clipPointsDefault(true)
drawToScale(true)
drawToScale(false)
maxPoints(500)
minPoints(21)
screenResolution(600)
--Operations which return default information:
adaptive()
clipPointsDefault()
drawToScale()
maxPoints()
minPoints()
screenResolution()

———-

References^{[1] nothing}

4